OpenStudy (anonymous):

estimate 812-349 using front-end-estimation show work

OpenStudy (akshay_budhkar):

eh? use the calci?

OpenStudy (karatechopper):

idk what front end elstimation is but i know reg estimation

OpenStudy (karatechopper):

800-350

OpenStudy (anonymous):

812 = 800 + 12 349 = 350 - 1 800 - 350 +12 + 1 = 450 + 12 + 1 = 463

OpenStudy (akshay_budhkar):

omg i made a fool of myself lol.. never heard of such a method o_O

OpenStudy (anonymous):

i never heard of it either

OpenStudy (anonymous):

me etheir

OpenStudy (karatechopper):

akshay same here

OpenStudy (akshay_budhkar):

robtobey is typing a reply… maybe he knows it

OpenStudy (karatechopper):

u mean ankonsha then it just became robtobey

OpenStudy (anonymous):

A hand calculator is the way to go.

OpenStudy (anonymous):

this is kind ah confusing

OpenStudy (akshay_budhkar):

? that took a while @robtobey

OpenStudy (anonymous):

so wht answer

OpenStudy (akshay_budhkar):

i will go with 463

OpenStudy (anonymous):

are u sure it is 146

OpenStudy (akshay_budhkar):

?

OpenStudy (anonymous):

wht is the answer

OpenStudy (akshay_budhkar):

463

OpenStudy (anonymous):

ok